# Time-dependent Hamiltonians and Geometry of Operators Generated by Them

### v1

@@ IV.3 The complexity geometry
where $\lambda_{1}^{2}$ and $\lambda_{2}^{2}$ are two constants that incorporate the numerical factors from the first and second terms, respectively, and we have also renamed $\lambda_{12}^{2}=\lambda_{1}^{2}+\lambda_{2}^{2}$ in the second expression. Clearly, the role of the last term in the second expression above (which is a part of the contribution of the non-diagonal matrix elements of the Hamiltonian towards the norm) is to reduce the total cost. An alternative expression for the cost function can be obtained in terms of the expectation value of the Hamiltonian in the state $\big{|}\phi_{0}(t)\big{>}$ (and the frequency) by noting that the function $f(t)$ is related to it by Eq. (40) so that we have,
@@
After deriving this expression for the cost we need to show that in an appropriate limit this reduces to the expression for the HO obtained in YK and discussed in sec- III.1. To check this, we notice that in the limit $\Omega(t)=\omega=\text{constant}$, it is possible to directly integrate Eq. (39) to obtain a first integral of motion (this can be directly seen from the general formula derived in Eq. (43) by putting $\dot{\Omega}=0$). With the integration constant written as $2\omega$ (with $\omega>0$) we have LR
@@
$\Big{(}\mathcal{C}\big{(}\mathbf{H}_{\Omega}(t)\big{)}\Big{)}^{2}=\bigg{[}\int_{0}^{T}\sqrt{\lambda_{12}^{2}f^{2}(t)-4\lambda_{2}^{2}\Omega^{2}(t)}~{}\text{d}t\bigg{]}^{2}\\ ~{}\leq~{}T\bigg{[}\int_{0}^{T}\bigg{(}\lambda_{12}^{2}f^{2}(t)-4\lambda_{2}^{2}\Omega^{2}(t)\bigg{)}~{}\text{d}t\bigg{]}~{}\leq~{}\lambda_{12}^{2}T\int_{0}^{T}~{}f^{2}(t)~{}\text{d}t=\lambda_{12}^{2}T^{2}\overline{f^{2}(t)}~{},$ (60)
with $\overline{f^{2}}$ denoting the time average of the function $f(t)$. Combining the final and the initial expressions above, we have, for an arbitrary point on the unitary evolution curve generated by a OTF, the following upper bound on the total cost
$\mathcal{C}\big{(}\mathbf{H}_{\Omega}(t)\big{)}~{}\leq~{}\lambda_{12}T\sqrt{\overline{f^{2}(t)}}~{}=~{}\lambda_{12}T\sqrt{\overline{\Big{(}\big{<}\phi_{0}(t)\big{|}\mathbf{H}_{\Omega}(t)\big{|}\phi_{0}(t)\big{>}\Big{)}^{2}}}~{}.$ (61)
@@
$T~{}\geq~{}\frac{\mathcal{C}\big{(}\mathbf{H}_{\Omega}(t)\big{)}}{\lambda_{12}\sqrt{\overline{f^{2}(t)}}}~{}=~{}\frac{\mathcal{C}\big{(}\mathbf{H}_{\Omega}(t)\big{)}}{\lambda_{12}\sqrt{\overline{\Big{(}\big{<}\phi_{0}(t)\big{|}\mathbf{H}_{\Omega}(t)\big{|}\phi_{0}(t)\big{>}\Big{)}^{2}}}}~{}.$ (62)
Inequality such as this often appears in the context of the quantum speed limit, which is usually defined as the minimum amount of time required to prepare a quantum state with unit fidelity. See, e.g. Bukov , where such a geometric bound for the quantum speed limit has been conjectured.
To illustrate the results presented so far in this section, we now consider some OTF models that are commonly studied in the literature. Later in Appendix. A, we discuss one further example of a time-dependent protocol, namely a quantum quench scenario, which is one of the most popular protocols for studying non-equilibrium dynamics in quantum systems, and find out the behaviour of different physical quantities in adiabatic as well as fast quench regimes.
@@ IV.4 Example-1 : Instantaneous frequency with non-vanishing Schwarzian derivative
$\mathcal{C}\big{(}\mathbf{H}_{\Omega}(t_{f})\big{)}=\int_{0}^{t_{f}}~{}F_{r}\big{(}\mathbf{H}_{\Omega}(t)\big{)}~{}\text{d}t=\frac{\lambda_{3}}{b_{2}}\ln\big{[}1+\frac{b_{2}}{b_{1}}t_{f}\big{]}~{}.$ (66)
@@ V.1 Computation of the norm
where we have defined three time-dependent functions as
@@
The infinite summation above is once again an example of the generalised Epstein-Hurwitz series, and we call its numerical value (including the factor of $\lambda_{0}^{2}$) to be $\lambda_{2}^{2}$, so that the final expression for the norm is given by
@@
$F^{2}_{r}\big{(}\mathbf{H}_{m_{1}\Omega_{1}}(t)\big{)}=\lambda_{1}^{2}f^{2}_{3}(t)+\lambda_{2}^{2}\Big{(}f^{2}_{3}(t)-4A_{0}(t)B_{0}(t)+C_{0}^{2}(t)\Big{)}~{},$ (84)
@@
V.2 Example - Complexity geometry associated with the Caldirola–Kanai Hamiltonian
V.3 $su(1,1)$ algebra and the time evolution operator
From now on, we exclusively consider the time-evolution operator and bi-invariant cost of preparing an arbitrary operator on the unitary time evolution curve. This is the solution given in Eq. (3) of the corresponding Schrodinger equation (1), with the path order replaced by the usual time ordering. In many cases, a compact expression for the time evolution operator ($\mathcal{U}(t)$) can be derived with the help of the Lie group theory Dattoli:1988zza ; CFL . Below, we see that the Hamiltonian of OTMF is one such example, and thus, it is possible to obtain a compact expression for the corresponding time evolution operator and study various aspects of its associated bi-invariant cost.
@@ V.3 $su(1,1)$ algebra and the time evolution operator
These relations can also be established by using a $2\times 2$ matrix representation of the corresponding group element. [^12] See Appendix B for a brief derivation of these decomposition formulas using the matrix representations. Given the time dependence of the mass and angular frequency of the OTMF, we can determine the functions $c_{i}$ by using the following relations CFL
@@
$\frac{d^{2}x_{c}(t)}{dt^{2}}+\xi(t)\frac{dx_{c}(t)}{dt}+\Omega_{1}^{2}(t)x_{c}^{2}(t)=0~{},\quad\text{with}\quad\xi=\frac{d\ln m_{1}(t)}{dt}~{},$ (95)
@@
V.4 Equivalence between OTF and OTMF under time reparametrisation
@@
and vice versa. Here, $c_{3}$ is an integration constant. In the following, we assume that the integral above can be performed for the range of the parameter $t$ of our interest and the resulting expression for $\tau$ can be inverted to obtain the function $t(\tau)$. This is certainly the case for the examples we have discussed in this paper. For example, for the CK oscillator considered in the sec - V.2, these two parameters are related by the formula
@@
V.5 Equivalence between the cost functional of the time evolution operators for the OTF and OTMF
@@
In view of this relation between the total costs, here we notice the following points:
@@
2. It is also important to notice that the relation between the total costs we have established is not an artifact of the unitary invariance of the cost function. The cost functionals are equivalent even if the two Hamiltonians are not related by any time-independent unitary transformation. In fact, the two Hamiltonians in question are not related by any time-independent unitary transformation. The equivalence of costs directly follows from the definitions of the time evolution operator itself. During the calculation of the cost of creating a certain operator generated by a quantum mechanical Hamiltonian we can always perform such a reparametrisation of $t$, however, unlike the “special” transformation given in Eq. (96), corresponding norms of the generators do not have any physical meaning as the Hamiltonian operator of certain quantum mechanical system as such. Furthermore, it will be interesting to find out whether such mapping of costs can be found for other pairs of generators, i.e. whether it is possible to find out other pairs of ‘cost-equivalent’ generators such that other quantities (e.g., wavefunction or the time evolution operator) also map to each other under the same reparametrisation which makes costs equivalent to each other.
3. Before moving on, we observe two more points related to this mapping between two systems. Firstly, the norms of the two Hamiltonians do not map to each other under the said transformation; it is only their integrals, i.e. the cost functional or the total cost of creating a certain operator which map to each other. The norms of the Hamiltonians themselves can be quite different from each other. In this context we also want to point out the importance of the relation in Eq. (100) : The presence of the extra factor of $B_{1}$ in this relation cancels out the $1/B_{1}$ factor from the time reparametrisation in the total cost so that the equivalence between the cost functional holds.
The next point to notice is that the Hamiltonians $\mathbf{H}_{m_{1}\Omega_{1}}(t)$ in Eq. (35) and $\mathbf{H}_{\Omega}(\tau)$ in Eq. (73) can be formally connected to each other through suitable unitary transformations (see e.g. Mostafazadeh:1998hq where such transformations were derived). However, these unitary transformations are time-dependent, and therefore, the bi-invariant costs of these two Hamiltonians are not equivalent.
V.6 Cost and complexity along the time evolution curve generated by time-dependent Hamiltonians
@@
Before concluding this section we note the following: Given a particular operator at an arbitrary point on the time evolution curve, it would be interesting to define a efficiency associated with the evolution, following, e.g., Anandan , as the ratio of the minimum (geodesic) distance computed along the path connecting the identity and the final operator (or the initial and the final states), and the length of the actual path generated by the time-dependent Hamiltonian under consideration. The difference between the efficiency introduced in Anandan and our proposal is that the former uses the Fubini-Study metric defined on the space pure quantum state to measure the distance, [^14] See Cafaro for a recent work that uses this definition to measure the efficiency of the analog quantum search algorithms. while our definition uses the bi-invariant CG to measure the same. It will be interesting to study this quantity further in a future work.
VI Summary and discussions
In this paper we have discussed a method of obtaining the bi-invariant CG of the operator generated by the time-dependent Hamiltonian of a quantum mechanical system. The procedure we have followed makes use of a regularisation scheme in which the norm of an infinite dimensional Hamiltonian is obtained by first imposing a condition on its trace and subsequently using the resultant mean value ($a$) in the calculation of the cost function. For time-independent Hamiltonians (which we had considered in section III), this mean value is also a constant, and therefore, the cost itself is independent of time. However, for explicitly time-dependent Hamiltonians considered in later sections, this mean value is itself a function of time, and as a result, the computation of the regularised cost is more involved. Another feature of the calculation where the computation of the cost of the two types of systems differs is that, for time-independent Hamiltonians, one only needs to know the energy eigenvalues to calculate the cost, whereas since a time-dependent Hamiltonian has no stationary eigenstates, both diagonal as well as non-diagonal matrix elements of this Hamiltonian in a suitable basis are required for the computation of the cost. To this end, we have used the well known Lewis-Riesenfeld method of time-dependent invariant operators since the eigenfunctions of this operator can be directly used to efficiently find out both the diagonal and non-diagonal matrix elements of the Hamiltonian in question, and hence the norm can be calculated in terms of these matrix elements. This is the strategy we have followed in section IV and V to obtain the regularised norm of time-dependent Hamiltonians, where the generic properties of the obtained results are also discussed.
As emphasised in the main text, we again want to stress the importance of looking at the familiar problem of computational complexity from a slightly different viewpoint: rather than finding a minimum of all unitary curves connecting the two given operators (which is the standard framework of Nielsen’s approach), calculating the total cost of evolution along a specified path on a CG is also important, since the time-dependent Hamiltonian of a quantum mechanical system may itself generate such unitary curve. The assumption of bi-invariant geometry and the regularisation procedure used to find out the norm for an infinite dimensional Hilbert space are two key ingredients of our analysis, and both have strong effects on the final form of the cost function we have derived.
Among various examples of explicitly time-dependent systems, in this paper, we have focused on two particular Hamiltonians, which are widely studied in the literature due to their importance in modelling various physical phenomena. These are, respectively, an oscillator of unit mass whose frequency of oscillation explicitly depends on time and an oscillator whose both mass and frequency change with time. For both these cases, using the Lewis-Riesenfeld method, we have obtained compact analytical expressions for the cost of the time evolution operator generated by these two Hamiltonians. We have also obtained a formal equivalence between the two separate cases considered so that expression for various quantities of two systems can be mapped to one another by a time reparametrisation, and demonstrated the implications of this duality in terms of the total bi-invariant cost. Finally, another important example of time-dependent quantum system, namely a smooth quench of the Hamiltonian parameters is considered in Appendix A. For a time-dependent frequency oscillator model, using the exact analytical expressions available in the literature, we have studied information theoretic quantities, e.g., change of the Shannon entropy, and compared them with the cost of the time-evolution operator derived here. In both cases we have shown that different quench regimes and their associated characteristics can be effectively probed by these quantifiers.
Before concluding, we mention a few possible future directions. Firstly, in this paper we have mainly considered the generalised oscillator and other Hamiltonians which can be connected to it by unitary transformations. Extending our analysis to systems beyond this class, specifically systems that are classically chaotic, can be one important direction that can be pursued. Secondly, the bi-invariant CG we have discussed in this paper is invariant under unitary transformations between the generators which are independent of the parameter describing the unitary curve, e.g., as we have discussed in V.1, the unitary transformation $U_{1}$ in (72) connecting the Hamiltonians of a generalised oscillator and an OTMF is time-independent, hence, the cost is the same for both the Hamiltonians. However, if we consider two Hamiltonians which are related through a time-dependent unitary transformation, the CG of Eq. (11) is no longer invariant under such transformation (this is the reason why the costs associated with OTF and OTMF are not the same). Determining an invariant CG generated by Hamiltonians related through general time-dependent unitary transformations is an interesting problem, and we hope to come back to it in a future work.
Acknowledgments
We sincerely thank Hugo A. Camargo, Yichao Fu, Viktor Jahnke, Keun-Young Kim, and Run-Qiu Yang for useful discussions on this and other related works, as well as comments on a draft version of the manuscript. We would also like to thank the organisers of the workshop “Student Talks in Trending Topics in Theory” (ST4 -2023), during which parts of the present work were carried out. The work of Kunal Pal is supported by the National Research Foundation of Korea under Grants No. 2017R1A2A2A05001422 and No. 2020R1A2C2008103.
Appendix A Quantum quenches - an exact example of time-dependent scenarios

### v2

@@ IV.3 The complexity geometry
where $\lambda_{1}^{2}$ and $\lambda_{2}^{2}$ are two constants that incorporate the numerical factors from the first and second terms, respectively, and we have also renamed $\lambda_{12}^{2}=\lambda_{1}^{2}+\lambda_{2}^{2}$ in the second expression. This constitutes one of the main results of the present paper. Clearly, for a normal oscillator, for which the frequency is always a real function of time, the role of the last term in the second expression above (which is a part of the contribution of the non-diagonal matrix elements of the Hamiltonian towards the norm) is to reduce the total cost. An alternative expression for the cost function can be obtained in terms of the expectation value of the Hamiltonian in the state $\big{|}\phi_{0}(t)\big{>}$ (and the frequency) by noting that the function $f(t)$ is related to it by Eq. (30) so that we have,
@@
After deriving this expression for the cost we need to show that in an appropriate limit this reduces to the expression for the HO obtained in YK and discussed in section III.1. To check this, we notice that in the limit $\Omega(t)=\omega=\text{constant}$, it is possible to directly integrate Eq. (29) to obtain a first integral of motion (this can be directly seen from the general formula derived in Eq. (33) by putting $\dot{\Omega}=0$). With the integration constant written as $2\omega$ (with $\omega>0$) we have LR
@@
$\big{(}\mathcal{C}_{\Omega}(T)\big{)}^{2}=\bigg{[}\int_{0}^{T}\sqrt{\lambda_{12}^{2}f^{2}(t)-4\lambda_{2}^{2}\Omega^{2}(t)}~{}\text{d}t\bigg{]}^{2}\\ ~{}\leq~{}T\bigg{[}\int_{0}^{T}\bigg{(}\lambda_{12}^{2}f^{2}(t)-4\lambda_{2}^{2}\Omega^{2}(t)\bigg{)}~{}\text{d}t\bigg{]}~{}\leq~{}\lambda_{12}^{2}T\int_{0}^{T}~{}f^{2}(t)~{}\text{d}t=\lambda_{12}^{2}T^{2}\overline{f^{2}(t)}~{},$ (50)
with $\overline{f^{2}}$ denoting the time average of the function $f(t)$ up to time $T$. Combining the final and the initial expressions above, we have, for an arbitrary point on the unitary evolution curve generated by a OTF, the following upper bound on the total cost
$\mathcal{C}_{\Omega}(T)~{}\leq~{}\lambda_{12}T\sqrt{\overline{f^{2}(t)}}~{}=~{}\lambda_{12}T\sqrt{\overline{\Big{(}\big{<}\phi_{0}(t)\big{|}\mathbf{H}_{\Omega}(t)\big{|}\phi_{0}(t)\big{>}\Big{)}^{2}}}~{}.$ (51)
@@
$T~{}\geq~{}\frac{\mathcal{C}_{\Omega}(T)}{\lambda_{12}\sqrt{\overline{f^{2}(t)}}}~{}=~{}\frac{\mathcal{C}_{\Omega}(T)}{\lambda_{12}\sqrt{\overline{\Big{(}\big{<}\phi_{0}(t)\big{|}\mathbf{H}_{\Omega}(t)\big{|}\phi_{0}(t)\big{>}\Big{)}^{2}}}}~{}.$ (52)
Inequality such as this often appears in the context of the quantum speed limit Deffner:2017cxz , which is usually defined as the minimum amount of time required to prepare a quantum state with unit fidelity. See, e.g. Bukov , where such a geometric bound for the quantum speed limit in quantum many-body systems has been conjectured.
To illustrate the results presented so far in this section, we now consider some OTF models that are commonly studied in the literature. Later in section VI, we discuss one further example of a time-dependent protocol, namely a quantum quench scenario, which is one of the most popular protocols for studying non-equilibrium dynamics in quantum systems, and find out the behaviour of different physical quantities in adiabatic as well as fast quench regimes.
@@ IV.4 Example-1 : Instantaneous frequency with non-vanishing Schwarzian derivative
$\mathcal{C}_{\Omega}(t_{f})=\int_{0}^{t_{f}}~{}F_{r}\big{(}\mathbf{H}_{\Omega}(t)\big{)}~{}\text{d}t=\frac{\lambda_{3}}{b_{2}}\ln\big{[}1+\frac{b_{2}}{b_{1}}t_{f}\big{]}~{}.$ (56)
@@ V.1 Computation of the norm
where we have defined three time-dependent functions through the relations
@@
The infinite summation above is once again an example of the generalised Epstein-Hurwitz series, and we call its numerical value (including the factor of $\lambda_{0}^{2}$) to be $\lambda_{2}^{2}$, so that the expression for the norm is given by
@@
$\begin{split}F^{2}_{r}\big{(}\mathbf{H}_{m_{1}\Omega_{1}}(t)\big{)}=\lambda_{1}^{2}f^{2}_{3}(t)+\lambda_{2}^{2}\Big{(}f^{2}_{3}(t)-4A_{0}(t)B_{0}(t)+C_{0}^{2}(t)\Big{)}~{}\\ =\lambda_{12}^{2}f^{2}_{3}(t)-\lambda_{2}^{2}\Big{(}4A_{0}(t)B_{0}(t)-C_{0}^{2}(t)\Big{)}~{},\end{split}$ (74)
@@
In the second expression for the cost function above, we have written it in a form that is similar to the cost of OTF in Eq. (47). If we assume that the time-dependent coefficients in the Hamiltonian (63) are such that the frequency of OTMF is always a positive function of time, and $B_{0}(t)>0$, then the role of the second term is to reduce the cost function. In Appendix B we provide an example of computation of this cost function for given time-dependent profiles of mass of frequency of the oscillator.
In this case we can derive a bound on the total cost by following a similar procedure as in section IV.3 to be
$\mathcal{C}_{m_{1}\Omega_{1}}(T)~{}\leq~{}\lambda_{12}T\sqrt{\overline{f_{3}^{2}(t)}}~{},$ (76)
where $T$ is final time up to which the total cost has been calculated, and $\overline{f_{3}^{2}(t)}$ denotes time average of $f_{3}^{2}(t)$ up to the final time.
V.2 $su(1,1)$ algebra and the time evolution operator
We now explicitly consider the time-evolution operator associated with the OTMF, which is essentially the solution given in Eq. (3) of the corresponding Schrodinger equation (1), with the path order replaced by the usual time ordering. In many cases, a compact expression for the time evolution operator ($\mathcal{U}(t)$) can be derived with the help of the Lie group theory Dattoli:1988zza ; CFL . Below, we see that the Hamiltonian of OTMF is one such example, and thus, it is possible to obtain a compact expression for the corresponding time evolution operator and study various aspects of its associated bi-invariant cost.
@@ V.2 $su(1,1)$ algebra and the time evolution operator
These relations can also be established by using a $2\times 2$ matrix representation of the corresponding group element. [^10] See Appendix C for a brief derivation of these decomposition formulas using the matrix representations. Given the time dependence of the mass and angular frequency of the OTMF, we can determine the functions $c_{i}$ by using the following relations CFL
@@
$\frac{d^{2}x_{c}(t)}{dt^{2}}+\xi(t)\frac{dx_{c}(t)}{dt}+\Omega_{1}^{2}(t)x_{c}^{2}(t)=0~{},\quad\text{with}\quad\xi(t)=\frac{d\ln m_{1}(t)}{dt}~{},$ (83)
@@
V.3 Equivalence between OTF and OTMF under time reparametrisation
@@
and vice versa. Here, $c_{3}$ is an integration constant. In the following, we assume that the integral above can be performed for the range of the parameter $t$ of our interest and the resulting expression for $\tau$ can be inverted to obtain the function $t(\tau)$. This is certainly the case for the examples we have discuss in this paper. For example, for the Caldirola–Kanai oscillator considered in the Appendix. B, these two parameters are related by the formula
@@
V.4 Equivalence between the cost functional of the time evolution operators for the OTF and OTMF
@@
In view of this relation between the total costs, here we discuss some important points which should further clarify the meaning and the implications of this equivalence.
@@
2. It is also important to notice that the relation between the total costs we have established is not an artifact of the unitary invariance of the cost function. The total costs are equivalent even if the two Hamiltonians are not related by any time-independent unitary transformation. In fact, the two Hamiltonians in question (an OTF and OTMF) are not related by any time-independent unitary transformation. The equivalence of costs directly follows from the definitions of the time evolution operator itself. During the calculation of the total cost of creating a certain operator generated by a given Hamiltonian we can always perform such a reparametrisation of $t$, however, unlike the ‘special’ transformation given in Eq. (84), the transformed total norm usually does not have any physical meaning as the integral pf the norm of the Hamiltonian operator of certain other quantum mechanical system. Furthermore, it will be interesting to find out whether such mapping of costs can be found for other pairs of generators, i.e. whether it is possible to find out other pairs of ‘(total) cost-equivalent’ generators such that other quantities (e.g., wavefunction or the time evolution operator) also map to each other under the same reparametrisation which makes total costs equivalent to each other.
3. Before moving on, we observe two more points related to this mapping between two systems. Firstly, the norms of the two Hamiltonians do not map to each other under the said transformation; it is only their integrals, i.e. the cost functional or the total cost of creating a certain operator which map to each other. The norms of the Hamiltonians themselves can be quite different. In this context, we also want to point out the importance of the relation in Eq. (88) : The presence of the extra factor of $B_{1}$ in this relation cancels out the $1/B_{1}$ factor from the time reparametrisation in the total cost so that the equivalence between the cost functional holds. In fact, this is also the reason why the upper bounds on the total cost functions in eqs. (51) and (76) do not map to each other under time reparametrisation.
Secondly, the Hamiltonians $\mathbf{H}_{m_{1}\Omega_{1}}(t)$ in Eq. (25) and $\mathbf{H}_{\Omega}(\tau)$ in Eq. (63) can be formally connected to each other through suitable unitary transformations (see e.g. Mostafazadeh:1998hq where such transformations were derived). However, these unitary transformations are time-dependent, and therefore, the bi-invariant costs of these two Hamiltonians are not equivalent.
V.5 Cost and complexity along the time evolution curve generated by time-dependent Hamiltonians
@@
Before concluding this section we note the following: Given a particular operator at an arbitrary point on the time evolution curve, it would be interesting to define an efficiency associated with the evolution, following, e.g., Anandan , as the ratio of the minimum (geodesic) distance computed along the path connecting the identity and the final operator (or the initial and the final states), and the length of the actual path generated by the time-dependent Hamiltonian under consideration. The difference between the efficiency introduced in Anandan and our proposal is that the former uses the Fubini-Study metric defined on the space pure quantum state to measure the distance, [^12] See Cafaro for a recent work that uses this definition to measure the efficiency of the analog quantum search algorithms. while our definition uses the bi-invariant CG to measure the same. It will be interesting to study this quantity further in a future work.
VI Quantum quenches - an exact example of time-dependent protocols